OpenStudy (anonymous):

Which of the following is the best linear approximation for f(x) = cos(x) near x = pi/2 ?

OpenStudy (perl):

we can approximate y(x) with L(x) y(x) ~ L(x) , where L(x) = y (a) + y ' (a) ( x - a)

OpenStudy (perl):

where the approximation is very good as x -> a

OpenStudy (perl):

L(x) = cos(pi/2) + (-sin (pi/2)) * (x - pi/2)

OpenStudy (anonymous):

which is L(x) = 0 + (- 1) * (x - pi/2)

OpenStudy (anonymous):

-x + pi/2

OpenStudy (perl):

yes that looks correct